OpenStudy (anonymous):

Write the expression in terms of the logarithms of a, b, and c. ln((sqrt(ab^8))/c)

OpenStudy (anonymous):

\[\begin{align} \ln\left(\frac{\sqrt{ab^8}}{c}\right)&=\ln(\sqrt{ab^8})-\ln(c) \\ &=\ln[(ab^8)^{1/2}]-\ln(c) \\ &=\frac{1}{2}\ln(ab^8)-\ln(c) \\ &=\frac{1}{2}[\ln(a)+\ln(b^8)]-\ln(c) \\ &=\frac{1}{2}[\ln(a)+8\ln(b)]-\ln(c) \\ &=\frac{1}{2}\ln(a)+4\ln(b)-\ln(c) \end{align}\]

OpenStudy (anonymous):

These question usually involve the applicaiton of the three logarithm rules (as \(\ln(x)\) is just another logarithm!): \[\log_z\left(\frac{x}{y}\right)=\log_z(x)-\log_z(y) \\ \log_z\left({x}{y}\right)=\log_z(x)+\log_z(y) \\ \log_zx^y=y\log_zx\]